I need help with this equation please: Supposed to solve the equation and check for extraneous solutions.
3√(w+1)=6
Thanks

Answer by edjones(8007)

3sqrt(w+1)=6
sqrt(w+1)=2
w+1=4 Square each side.
w=3
.
Ed